### v2

# Dark Matter Substructure or Source Model Systematics? A Case Study of Cluster Lens Abell S1063

Abstract
Mapping the small-scale structure of the universe through gravitational lensing is a promising tool for probing the particle nature of dark matter. Curved Arc Basis (CAB) has been proposed as a local lensing formalism in galaxy clusters, with the potential to detect low-mass dark matter substructure. In this work, we analyze the cluster lens Abell S1063 in search of dark matter substructure with the CAB formalism, using multi-band imaging data from JWST. We use two different source modeling methods: shapelets and pixel-based source reconstruction based on Delaunay triangulation. We find that source modeling systematics from shapelets result in a disagreement between CAB parameters measured from different filters. Source modeling with Delaunay significantly alleviates this systematic, as seen in the improvement in agreement across filters. We also find that inadequate complexity in source modeling can result in convincing spurious detections of dark matter substructure from strong gravitational lenses, as seen by our measurement of a subhalo with shapelets, a spurious detection that is not reproduced with Delaunay source modeling. We demonstrate that multi-band analysis with different JWST filters is key for disentangling source and lens model systematics from dark matter substructure detections.
keywords:
gravitational lensing: strong – dark matter – galaxies: clusters1 Introduction
The standard model of cosmology (CDM) assumes the existence of Cold Dark Matter (CDM), which dominates the matter content of the universe and is essential to explaining an array of cosmological observables spanning cosmic epochs. These include observations of the cosmic microwave background (Planck Collaboration et al. 2020) and large-scale structure of the universe (Springel et al. 2006,Gil-Marín et al. 2016, Abbott et al. 2018, Troxel et al. 2018, Abolfathi et al. 2018, Adame et al. 2025). While it has been remarkably successful at large scales, the cold dark matter paradigm has yet to be tested on small scales. Simulations with CDM predict an abundance of low-mass structures called subhalos residing in dark matter halos around galaxies (de Lucia et al. 2004, Chan et al. 2015). At the same time, many alternative dark matter theories can reproduce our large-scale observations but predict distinct observational signatures at sub-galactic scales. These theories either modify the predicted density profiles and population statistics of subhalos, as seen in warm or self-interacting dark matter models (Bode et al. 2001, Tulin & Yu 2018), or predict entirely different forms of substructure, such as interference patterns arising in ultralight axion dark matter (Schive et al. 2014). To this end, probing the small-scale structure of the universe can be the key to understanding the physics of dark matter.
Strong gravitational lensing is sensitive to the line-of-sight mass distribution between the source and the observer (Mao & Schneider 1998), making it a direct probe of the dark matter substructure. As predicted by general relativity, light rays bend around massive objects, such as galaxies or galaxy clusters, leading to the appearance of multiple distorted images of a background source. The mass distribution of the lens is often parametrized by a smooth profile. A dark matter clump, either within the lens or along the line of sight, will cause a perturbation on top of this smooth lens model, allowing us to detect dark matter substructure gravitationally. To this day, a few subhalos have been detected or reported as candidate detections with this method in galaxy-galaxy strong lenses (Vegetti et al. 2010, Hezaveh et al. 2016, Şengül et al. 2022, Lange et al. 2025).
Cluster lenses are even more powerful than galaxies, as they often host rich collections of multiply imaged, highly magnified, distorted arcs. However, studies of dark matter substructure in cluster lenses have been hindered by the computational cost of analyzing pixel-level data from their extensive fields of view. To this end, typical strong lens modeling of clusters has involved reducing the rich pixel-level data to only the positional information of lensed images (e.g., Lagattuta et al. 2019). However, recent work has made significant progress on incorporating pixel-level information in global cluster models (Acebron et al. 2024, Broadhurst et al. 2025). As a result, cluster substructure studies have been limited to statistical constraints on subhalo population abundances (Natarajan et al. 2007, Natarajan et al. 2017). While such studies give us important insights into the abundance of dark matter substructure, probing the density profiles of individual subhalos with direct detection techniques allows us to go a step further. By comparing the measured concentration, ellipticity, and effective density slope (Şengül & Dvorkin 2022) of subhalos to results from simulations with different dark matter models, we can support or rule out dark matter particle candidates and place tighter constraints on the viable parameter spaces of models.
A promising approach for direct detection of cluster substructure is to employ a local lensing formalism. Curved Arc Basis (CAB) (Birrer 2021) was proposed as a local description of gravitational distortions based on the eigenvalues and eigenvectors of the local lensing Jacobian, the curvature of the tangential eigenvector, and an orientation of the curved arc. Şengül et al. (2023) showed with simulations that dark matter substructure along the line of sight of highly magnified lensed arcs in galaxy clusters can be detected and studied with the CAB formalism. They also had the first application of CAB to real data by modeling lensed images of the first cluster lens SMACS 0723 observed by the James Webb Space Telescope (JWST).
In this work, we exploit the power of multi-filter analysis of JWST data to study the source, lens, and foreground systematics involved in substructure searches with the CAB formalism. We describe the observations of Abell S1063 in Section 2. We detail our source and lens modeling methods, as well as substructure search in Section 3. We present our results of modeling a triply-lensed system in Abell S1063 with the CAB formalism and our substructure search in Section 4. We present our concluding thoughts in Section 5.
2 Abell S1063
Abell S1063 (also known as RXC-J2248-ID3) is a rich cluster lens at redshift . It was first identified by Abell et al. (1989) and was later observed as part of the Cluster Lensing And Supernova survey with Hubble (CLASH) program (Postman et al. 2012). Caminha et al. (2016) used the latter imaging data together with spectroscopic follow-up observations with the VIsible Multi-Object Spectrograph (VIMOS) and Multi Unit Spectroscopic Explorer (MUSE) on the Very Large Telescope (VLT) to identify 47 lensed images and construct a global mass model. Most recently, Beauchesne et al. (2024) combined these observations with X-ray emission from Chandra and velocity dispersion measurements of the member galaxies from MUSE to present an improved mass model.
2.1 Triply-imaged system 4
We use the detailed studies of Abell S1063 to carefully select a set of multiple images best suited for study with the CAB formalism. We choose a set of three images identified as system 4 by Caminha et al. (2016) for a number of reasons:
The images are clearly separated, indicating that they do not cross critical curves, where CAB formalism breaks down.
A set of three images mitigates lens modeling degeneracies.
The images are highly magnified, with a high signal-to-noise ratio, which is important for a detailed substructure search.
High-resolution JWST observations are available, and the images are visible across different filters. This allows for independent inference of the lens-mass distribution that should agree across filters.
The source has a complex, asymmetric morphology, making it more sensitive to perturbations in lensing due to substructure.
The local field around each image is free of line-of-sight foregrounds and nearby cluster member galaxies that often contaminate the lensed light of multiple images and can bias CAB parameters, complicating perturber searches.
The wider field of view of Abell S1063 is shown in Figure 1, where the positions of the three images of system 4 are indicated with red squares.
2.2 JWST Observations
We use James Webb Space Telescope observations111http://dx.doi.org/10.17909/mskr-6n50 of Abell S1063 proposed in the General Observing Program under Program 222https://www.stsci.edu/jwst/science-execution/program-information?id=1840 and taken on 24 September 2022 with the Near Infrared Camera (NIRCam). The data is available in three short-wavelength and three long-wavelength filters. We only analyze the short-wavelength filters (F115W, F150W, F200W) as they offer a sharper point spread function and twofold pixel resolution compared to the longer-wavelength filters. Both features are important to our analysis, as the substructure search relies on distinguishing pixel-level changes in the brightness of lensed images. The broadband filters F115W, F150W, and F200W are sensitive to a range of non-overlapping frequencies centered at pivot wavelengths of 1.154, 1.501, and 1.988 m, respectively. Thereby, each filter provides correlated but varying source morphologies lensed by the same mass distribution. Each filter’s exposure times and signal-to-noise ratios are reported in Table 1. The image cutouts of system 4 in the three filters are shown in Figure 2. The data used in our analysis is publicly available on the Mikulski Archive for Space Telescopes (MAST)333https://mast.stsci.edu/portal/Mashup/Clients/Mast/Portal.html.
| F115W | F150W | F200W | |
| Pivot Wavelength [m] | 1.154 | 1.501 | 1.988 |
| Exposure Time [s] | 1,245 | 1,031 | 1,245 |
| Signal-to-Noise Ratio | 110.1 | 130.9 | 140.6 |
| Pixel Size [arcseconds] | 0.031 | 0.031 | 0.031 |
We use the simulation tool WebbPSF444https://www.stsci.edu/jwst/science-planning/proposal-planning-toolbox/psf-simulation-tool to model the point spread function (PSF) of each NIRCam filter. Typically, the PSF can also be calculated from pointlike sources in the field of view. However, we turn to simulation software because field stars are scarce along the line of sight of Abell S1063, located at a galactic latitude of , in a direction away from the galactic plane.
3 Methods
Strong lensing analysis requires simultaneous modeling of the source light and lens-mass distribution, obtained by minimizing residuals between the observed data and model reconstructions. Our data (see Figure 2) consist of elliptical cutouts from pixel images of system 4 in Abell S1063 taken in three different filters. To model the cluster lensing, we use the local lensing formalism of CAB (Birrer 2021), which has been shown to work well in describing highly magnified, extended arcs in cluster lenses, as well as to be sensitive to dark matter substructure (Şengül et al. 2023). To capture the light distribution of the background galaxy, we use two alternative source modeling methods: parametric shapelets (Refregier 2003, Birrer et al. 2015, Birrer et al. 2021) and pixel-based Delaunay triangulation (Vegetti & Koopmans 2009, Şengül 2025). To perform statistical fitting and search for dark matter substructure, we use Bayesian nested sampling with dynesty (Speagle 2020).
3.1 Lens Modeling: CAB
3.1.1 Angular Deflection in CAB
The CAB formalism proposed by Birrer (2021) uses the eigenvalues and eigenvectors of the local lensing Jacobian to describe distortions of extended arcs. This local model is fully equivalent to a global Singular Isothermal Sphere with a mass sheet transformation (MST)555 MST is a transformation of the form , where and is the convergence field. Such transformations lead to a mass-sheet degeneracy as the shear signal is invariant under the addition of a constant surface mass density to the convergence field. and can be described by four local parameters: tangential stretch , radial stretch , orientation , and curvature . An illustration of the effect of a CAB distortion on a circularly symmetric source is shown in Figure 3. The deflection angles of a CAB transformation centered at can be expressed as:
where gets mapped to the center position by construction, and the angular deflection of a Singular Isothermal Sphere (SIS) centered at is given by:
The CAB formalism is especially relevant in cluster lenses, where the most massive component of the cluster dominates the strong gravitational lensing effect. Yang et al. (2020) have shown that, in some cases, lensed images in clusters may be well described by a constant convergence and shear, which is a special limit of the CAB as . Şengül et al. (2023) first applied CAB to data, modeling JWST images of the cluster lens SMACS 0723, and presented a novel method for detecting and studying low-mass subhalos. In particular, with the local lensing formalism in hand, one can detect subhalos located near highly magnified arcs by discerning their perturbative corrections to the main lensing by the cluster. In mock JWST-quality simulations of the SMACS 0723 system, the authors showed that the CAB formalism could detect perturbers down to masses of , allowing a measurement of their concentration, ellipticity, and redshift (Şengül et al. 2023).
3.1.2 Region of Validity
CAB is a local formalism and is only able to perfectly capture the lensing distortions for a global model that consists of a Singular Isothermal Sphere (SIS) and an allowed Mass Sheet Transform (MST), as given in Equation (1). Lenses that are more complex than SIS and MST will add corrections to CAB.
The ellipticity of the most massive cluster dark matter halo will induce derivatives along the tangential stretch of a curved arc. External shear will cause additional asymmetric distortions to the CAB lensing. These two effects can be accounted for by considering extensions of CAB incorporating tangential derivatives and shear, which add corrections to regular CAB away from the center of the image (Birrer 2021). However, the true lensing in a cluster is more complex than CAB and its extensions: the real mass distribution of clusters does not precisely follow these parametric profiles, an effect which can become important locally; at the same time, nearby galaxies in the field of view of the lensed images can induce additional lensing distortions to the images. The masses and concentrations of these galaxies are often poorly constrained, and incorporating them into the lens modeling can make the problem computationally intractable. Therefore, it has been emphasized that CAB is a local formalism that should be used cautiously (Birrer 2021). The region of local validity of a CAB approximation can easily be studied in simulations. However, the conclusions on the validity of CAB are heavily dependent on the specific configuration of our lensed images relative to the complex mass and light distribution of the cluster in question. In many cases, a detailed global model of a cluster lens may not be available. In this work, we take a general, empirical approach to illustrate the power of multi-band analysis with JWST in determining the region of CAB validity and disentangling lens model systematics. To this end, we perform our analysis in two different mask sizes applied to our JWST images. Masks are elliptical boundaries on the lens plane within which all the pixels are included in our analysis and outside of which all the pixels are excluded, illustrated with white dashed elliptical contours in Figure 2. If the local lensing formalism is a good approximation, assuming all other systematics are negligible, the parameters measured in different filters should be consistent. If the formalism fails due to inadequacies of CAB away from image centers, we should expect to see a tension between the same parameters measured in different filters, since different filters will have different spatial extents of source light. However, systematics can also arise from source modeling. To distinguish the two sources of systematics, we employ two different source modeling methods, as discussed in Section 3.2.
3.2 Source Modeling
Source modeling is an important part of strong gravitational lensing analysis, but it is especially important for CAB studies. JWST cluster data often involves more strongly magnified and higher-resolution lensed images than typical galaxy-galaxy strong lensing systems, thereby demanding a more sophisticated source model. In this work, we explore the strengths and limitations of two different source modeling methods: parametric shapelets and pixel-based Delaunay triangulation.
3.2.1 Shapelets
Shapelets (Refregier 2003, Birrer et al. 2015, Birrer et al. 2021) are an orthonormal set of weighted Hermite polynomials. A shapelet source model comprises a linear combination of the polynomial basis functions and is thereby fully specified by a list of shapelet coefficients, a reference scale, and a coordinate center. Shapelets are able to capture more complex light distributions than the common Sérsic (Sérsic 1963) profile of galaxies and have thus been commonmber of allowed basis functions through the relation . The size of the source is set by the reference scale parameter . The smallest scale resolved by order and reference scale is given by the relation . Finally, the light peak position of the source model is controlled by the shapelet center , which is a non-linear parameter in the model.
We determine the posterior distributions of and through nested sampling together with the lens model parameters, using uniform priors. At each sampling step, where the lens model is fixed, the light distribution on the lens plane is a simple linear combination of the shapelet basis function coefficients on the source plane. Thus, we can solve for the shapelet coefficients by a linear matrix inversion of an matrix, where the length of the data vector is the total number of pixels in our image cutouts. As the basis order is a hyperparameter, we perform a Bayesian Information Criterion (BIC) analysis to find the optimal with the lowest , where is the number of model parameters. This allows us to ensure that we use the optimal source complexity without overate BIC analyses for the shapelet basis order for each NIRCam filter, as the source model should naturally look different in different wavelength ranges in each filter, as well as for each lens model and mask size applied to image cutouts. All reported results use the shapelet that gives model fits with the lowest BIC score.
In our fiducial analysis with shapelets, we use a Cartesian basis, which is commonly used in lens modeling but is better suited to spherical source morphologies. However, if the source is highly elliptical, the implicit prior set by our choice of basis can bias the results. For this reason, we repeat our analysis using source reconstruction with an elliptical shapelet basis, which introduces two additional non-linear parameters. We set uninformative priors on these two parameters, allowing source ellipticity and position-angle of the semi-major axis to vary between and . Henceforth, refer to these two shapelet source reconstruction schemes with different bases as ‘Cartesian shapelets’ and ‘elliptical shapelets’. When a statement applies to both methods, we refer to them collectively as ‘shapelets’.
3.2.2 Delaunay Triangulation
We use a pixel-based source modeling method based on Delaunay triangulation as a non-parametric alternative to shapelets. The method was implemented in Şengül (2025) and is similar to the adaptive-grid technique introduced by Vegetti & Koopmans (2009) but differs in the algorithm for tesselation point selection. Most pixel-based modeling methods rely on various regularization methods to make pixel-level modeling more computationally tractable. However, regularization techniques can introduce new hyperparameters and modeling choices, such as demanding certain levels of smoothness in the source that may not always improve the accuracy of the source model.
In this work, we use the method described in Şengül (2025) to carefully select tesselation pixels in a way that makes pixel-based modeling tractable without demanding source regularization. The algorithm relies on selecting more tesselation points in the regions of the image plane where the light morphology is more complex. We use the gradient of the local surface brightness as an indicator of source complexity. We calculate the normalized Hessian at each pixel, where is the pixel brightness value, and choose candidate tesselation points, where is the total number of unmasked pixels in our data. If neighboring pixels are chosen as tesselation points, we remove one of each pair of pixels to prevent degeneracies in the source model reconstruction. At each sampling step, where the lens model is fixed, the tesselation points are mapped onto corresponding source plane points through backward ray-tracing. The list of light amplitudes at these points are calculated through a non-negative least squares optimization, as negative source model amplitudes are not physical. Our optimization is implemented using the scipy.optimize.nnls package, which is based on the method developed by Bro & De Jong (1997).
The list of vertex amplitudes fully determines our source model, as the source light can be reconstructed using Delaunay triangulation. On the source plane, we calculate the Delaunay triangulation of , dividing the plane into triangles. We calculate surface brightness values within each triangle by interpolating between the values at the vertices of the triangles:
where are the coefficients of the amplitudes at the three vertices , and
Any point that does not fall within a Delaunay triangle is assigned the surface brightness of the nearest vertex.
3.3 Bayesian Nested Sampling
We determine the source and lens models that best fit our observed data using a Bayesian framework. We independently analyze the images taken in three different NIRCam filters described in Section 2.2. For each filter, our data vector is an array of pixels in the image cutouts of System 4 shown in individual rows of Figure 2. Our model parameter vector consists of either only lens or both lens and source parameters, depending on our source modeling method,
since Cartesian and elliptical shapelets involve three or five non-linear source model parameters, respectively. Our model reconstruction is given by
where the lensing operator acts on the source light model to produce the lensed images, which are convolved with the PSF and then pixelized to an appropriate JWST-resolution grid, together denoted by operator, to give realistic image reconstructions. The source model is found through either a shapelets or a Delaunay triangulation reconstruction detailed in Section 3.2. The lensing operator is fully determined by the angular deflections of the CAB model given by Equations (1) and (2). The point spread function operation will depend on the NIRCam filter and determines the resolution of our observed lensed images in each filter. We implement all modeling in this work using lenstronomy (Birrer & Amara 2018, Birrer et al. 2021), a publicly available strong gravitational lensing package for modeling and simulations.
The likelihood of observing data given a set of lens model parameters is given by
where is the covariance matrix of the pixel errors, which combine the Poisson noise and instrumental read noise at each pixel in quadrature. The above expression assumes that the pixel errors are Gaussian. However, any systematics in the model can lead to deviations from a Gaussian distribution in the residuals (), potentially biasing posterior distributions of the parameters. Hence, a multi-filter analysis is crucial for illuminating any systematics and testing the robustness of our model.
The posterior distribution of model parameters is obtained by weighting prior distributions by the likelihoods,
In this work, we always assume uniform priors on all model parameters, listed in Table 2. We use dynesty (Speagle 2020) to implement this Bayesian analysis framework.
| Parameter Name | Units | Priors |
| [1.0, 7.0] | ||
| arcsec-1 | [, 0.5] | |
| [0.0, 1.0] | ||
| [0.0, 7] | ||
| [, 0.0] | ||
| arcsec-1 | [, 1.0] | |
| [, 0.5] | ||
| arcsec | [, 0.2] | |
| arcsec | [, 0.2] | |
| [0.0, 7.0] | ||
| [0.0, 7.0] | ||
| arcsec-1 | [, 1.0] | |
| [, 0.5] | ||
| arcsec | [, 0.2] | |
| arcsec | [, 0.2] | |
| arcsec | [0, 0.5] | |
| arcsec | [, 0.2] | |
| arcsec | [, 0.2] |
3.4 Substructure Search
We search for dark matter substructure following the method proposed in Şengül et al. (2023), by exploiting the proximity of potential subhalos to highly magnified regions of the lensed images. If a substructure is present along the line of sight of one of the images, it will cause a perturbation to the gravitational lensing sourced by the cluster, thereby referred to as a perturber. This signal would be picked up by a lens model with a dark matter subhalo on top of the CAB model. Şengül et al. (2023) showed in JWST-quality simulations that the CAB can be used to detect subhalos down to masses of and to measure their concentration, ellipticity, and redshift parameters.
We model the mass distribution of a dark matter substructure as a Navarro-Frenk-White profile (Navarro et al. 1996, Navarro et al. 1997), given by
which we parametrize in terms of , the mass enclosed within the virial radius , the radius at which the subhalo density is 200 times the critical density of the universe, and concentration , defined by the ratio . We treat the mass, concentration, and position of the perturber as free parameters in our lens model. To this end, we are agnostic to the nature of the substructure and do not assume, for example, that the subhalo follows a CDM mass-concentration relation. We place uniform priors on each subhalo parameter listed in Table 3 and search for a subhalo within from the brightest image pixel. We constrain the subhalo search to this region since the lensing effect of the perturbers is most discernible in regions with large gradients in the light distribution.
| Parameter | Initial | Wide | Narrow | Units |
| arcsec ["] | ||||
| arcsec ["] |
3.4.1 Source Regularization
Pixel-based source modeling methods commonly use regularization of the source light distribution (Warren & Dye 2003, Koopmans 2005, Suyu et al. 2006), where smoothness conditions are imposed on the source light to speed up the lens inversion. To investigate the effect of regularization on substructure detection, we extend our Delaunay triangulation algorithm to include gradient regularization introduced in Warren & Dye (2003). The modified merit function is given by:
where is the gradient regularization term and the is the regularization strength parameter. Hence, the linear source inversion problem becomes:
where is the regularization matrix. On a rectangular grid, typical gradient regularization penalizes differences in counts between neighboring pixels. In this work, we implement gradient regularization on a triangular grid by minimizing the differences between neighboring node values in a Delaunay mesh. The regularization term is a sum over all Delaunay triangles of the local gradient at the triangle, weighted by the areas of the triangles:
The regularization matrix is the matrix of all pairwise gradient dot products of :
As described in Section 3.2.2, we perform the regularized source inversion (Eq. 11) at each sampling step using a non-negative least squares optimization. We fit for the regularization parameter together with the lens model parameters in our Bayesian nested sampling framework.
4 Results
In this section, we present our results of modeling system 4 in Abell S1063 with CAB and performing a dark matter substructure search.
We first fit the lensed images with a smooth CAB model, using shapelet and Delaunay source modeling methods, and find improved agreement across filters with Delaunay, an indication of reduced systematics. Interestingly, our substructure search recovers a subhalo with shapelets but not with Delaunay, illustrating that insufficient source model complexity and inherent source regularization can lead to spurious substructure detections. We highlight the power of multi-band analysis in distinguishing dark matter substructure from source and lens model systematics.
4.1 Smooth CAB
We show the best-fit model reconstructions and residuals of a smooth CAB model obtained with Cartesian shapelets and Delaunay source modeling methods in Figures 4, 5 and 6, for filters F115W, F150W, and F200W, respectively. The Cartesian shapelet fits shown use the shapelet order parameters optimized for each filter. Interestingly, the best-fit source models reconstructed using the two methods vary considerably. As perhaps expected from a pixel-based method, Delaunay can reproduce more detail in the complex source light distribution composed of two bright peaks connected with a thin twisted arc. The residuals are lower with Delaunay than with Cartesian shapelet source modeling, although it is difficult to fairly compare the two fits given the different numbers of degrees of freedom involved in the two source modeling methods.
4.1.1 Analysis with Large Masks
We first analyze the images in three JWST filters using Large Masks, as shown in Figure 2. The posterior distributions of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling methods are shown in Appendix A. Each JWST filter offers an independent realization of the PSF and a different source morphology, but probes the same line-of-sight mass distribution. Therefore, the CAB parameters should agree across different JWST filters, and any tension between lens parameters must arise from systematics in our source or lens modeling. To quantify this tension, we compute for each parameter, where and denote the mean and standard deviation of a parameter measured with the th filter. Our results from both Cartesian shapelet and Delaunay source modeling methods show a significant average tension between lens parameters from different JWST filters. As discussed in Section 3.1.2, lens modeling systematics might arise from the failure of the local CAB approximation beyond its region of validity. This motivates us to shrink our mask sizes and consider smaller elliptical masks where we expect the CAB approximation to work better.
4.1.2 Analysis with Small Masks
We repeat the multi-filter analysis with a smooth CAB model using Small Masks (see Figure 2). We show the posterior distributions of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling in Appendix B. Reducing mask sizes leads to better agreement between filters with both source modeling methods, although a more significant improvement is seen with Delaunay. This shows that lens modeling systematics played an important role with Large Masks. Hence, multi-band analysis can be critical in determining the region of validity for a local lensing formalism.
Interestingly, Delaunay shows significantly better multi-filter agreement, with average tension between parameters at as compared to with Cartesian shapelets. Increasing the number of tesselation points to (see Section 3.2.2) in Delaunay source modeling further alleviates the tension to , a improvement from Cartesian shapelets. However, the average CT scores still point to considerable tension between parameters. Thus, systematic errors dominate the pixel errors, violating the assumption that pixel errors are Gaussian in Equation (7) and thereby biasing the posteriors. This motivates us to boost our JWST errros until the residuals from modeling failure become small compared to the boosted errors, burying systematics under Gaussian noise.
4.1.3 Artificially Boosted Noise
The tension between filters in our analysis with Small Masks indicates that our modeling systematics cannot fully handle the data quality of the JWST system 4 observations. To prevent systematics from biasing our measured CAB parameters, we boost the JWST error map by a factor of at each pixel. Such boosting of pixel errors was employed by Acebron et al. (2024) to limit the impact of imperfectly modeled areas on the lens model. We show the resulting posterior distributions from Cartesian shapelet and Delaunay source modeling in Figures 7 and 8, respectively, and compare the means and errors of each parameter in Figure 9. We find that several parameters measured by Cartesian shapelets are still in tension, with an average between different JWST filters. In contrast, source modeling with Delaunay results in a reduction in the average CT score to , bringing all parameters within a agreement between filters. We note that a two-fold boost in pixel errors leads to a factor of increase in the uncertainties of lens parameters. While error boosting does not solve the root cause of systematics, it allows us to adjust the data quality to a level that our systematics can successfully handle. With boosted errors, Delaunay performs significantly better than Cartesian shapelets, measuring consistent parameters. To reach the same level of consistency and thereby reliability with Cartesian shapelets, one would need to further boost pixel errors, reducing the quality of the data and the amount of information we can extract from it.
4.2 Substructure search
We perform a substructure search using a smooth CAB model and both shapelet and Delaunay source modeling methods. We parametrize the dark matter substructure with an NFW profile and sample for its mass, concentration, and 2D position on the lens plane, assuming a lens redshift for the subhalo (see Section 3.4). In our preliminary substructure search with Large Masks and Cartesian shapelet source modeling, we used wide uniform priors on all subhalo parameters, listed in Table 3 as Initial Priors. We limited the initial subhalo positional priors to arcseconds, where our coordinates are centered on the brightest image pixels, as proximity to bright regions is important for distinguishing the second-order lensing perturbation caused by subhalos. We searched for substructure in all three images of Abell S1063 system 4 and detected a subhalo in Image #2 with Cartesian shapelets. In our subsequent multi-filter analysis, we used tighter priors on the subhalo position centered at the preliminary detection location, referred to as Wide and Narrow Priors, listed in Table 3 and illustrated in Figure 10.
We find that source modeling systematics with shapelets can lead to convincing spurious detections of dark matter substructure. In particular, our substructure search with Cartesian shapelet source modeling results in a detection of a subhalo located at consistent positions in all three filters. And yet, the recovered subhalo concentration differs significantly across the three filters, hinting that the subhalo detection may be spurious. Indeed, these results are not reproduced with Delaunay source modeling, which we find is more robust against spurious substructure detection arising from the inadequate complexity of the source model.
4.2.1 Cartesian shapelets
We present the posterior distributions of the subhalo parameters recovered using Cartesian shapelet source modeling with Large Masks and Small Masks in Figures 11 and 12, respectively. We find that all NFW subhalo parameters are consistent at least between two different JWST filters. However, the concentration parameter measured in F150W consistently hits an upper prior of and is in tension with the posteriors of the F115W and F200W filters.
As discussed in Section 3.2, the complexity of the shapelet source model is controlled by the hyperparameter which should be optimized separately for the smooth and substructure lens models. We show the BIC curves of both models fitted to F115W, F150W, and F200W filters in Appendix C and compare the performance of smooth and substructure models at the optimal of each model. The BIC statistic consistently favors the subhalo model over a smooth CAB model in both F115W and F200 filters. On the other hand, we don’t see a significant difference in the BICs of smooth and substructure models for the F150W filter. It is possible to have a real substructure detection that is not reproduced in every filter if the SNR of those filters is insufficient to detect substructure (Lange et al. 2025). However, our lowest statistical significance for the subhalo is in the filter with an intermediate level of SNR (see Table 1), suggesting that it may be spurious.
We also show the posterior distribution of the subhalo parameters for a range of shapelet order parameters around the optimal in Appendix C. The subhalo parameters are generally consistent across a broad range of shapelet order parameters. This suggests that even the most complex Cartesian shapelet source models may suffer from inadequate source model complexity or other systematics, leading to a preference for a substructure model. However, in both F115W and F200W filters, the BIC between smooth and perturber models tends to decrease with increasing , another hint that the subhalo detection might be spurious.
We illustrate the detailed statistics of the subhalo detection with Cartesian shapelets in terms of the BIC and the logarithm of the Bayes factor in Figures 13 and 14, respectively. We repeated our substructure analysis using Small Masks and Large Masks and Wide and Narrow Priors imposed on the position of the subhalo in each of the three filters. With Large Masks and Narrow Priors, we consistently detect a subhalo with decisive statistics of and . Shrinking the positional priors of the subhalo to the region enclosed by Narrow Priors leaves F115W and F200W results unaffected but reduces the statistical significance of the subhalo detection in F150W. This is because the positions recovered with Wide Priors in the intermediate filter disagree with the other two filters, and imposing the subhalo to be near the position detected by the other two filters reduces its statistical significance. Interestingly, switching from Large to Small Masks generally reduces both the and of the subhalo detection in every filter. This suggests that lens modeling systematics arising from the failure of CAB formalism in larger masks also contribute to spurious subhalo detection.
Finally, we repeat our substructure search with boosted JWST errors (see Section 4.1.3), using Small Masks and Narrow Priors on the subhalo position. Notably, none of the filters continue to detect a subhalo with positive s. This provides further evidence that the detection of the subhalo was spurious. At the same time, the finding demonstrates the power of multi-band analysis. Looking for consistency in the lens parameters across filters allows us to decide whether we can trust our models. When tension across filters appear, we can use multi-band analysis to adjust the data quality to a level where systematics are drowned by boosted errors. To this end, we can prevent systematic effects from being mistaken for substructure detections.
4.2.2 Delaunay
We show the results of our substructure analysis relying on Delaunay source modeling in Figures 15 and 16. Interestingly, we do not detect a subhalo in any of the three filters, as and consistently favor a smooth model over added substructure in both Large and Small Masks and when using either Narrow or Wide Priors on the subhalo position. This is consistent with our null detection of a subhalo using Cartesian shapelets with boosted errors. As a result, we see that Delaunay is more robust against spurious subhalo detections. This can be explained by the fact that Delaunay source modeling suffers from lower systematics, as seen by better agreement across filters when using Delaunay instead of shapelet source modeling.
4.2.3 Elliptical Shapelets
We repeat our analysis with shapelets using an elliptical basis for source reconstruction. We show the resulting posterior distributions and subhalo detection statistics in Appendix D. As seen with Cartesian shapelets and Delaunay source modeling, the average CT score in Small Masks () is lower than that of Large Masks (), supporting our conclusion that a growing failure of the local lens model leads to higher systematics in increasing mask sizes. However, switching to an elliptical shapelet basis does not alleviate tension in CAB parameters across filters, as CT scores are comparable between Cartesian and elliptical shapelets. In both Large and Small Masks, a substructure model is favored over a smooth CAB model in only one filter (F200W), with and , respectively. Interestingly, in most filters and masks, the statistical significance of the subhalo detection is weaker as compared to Cartesian shapelets. This suggests that the spurious subhalo is statistically favored partially because of the failure of Cartesian shapelets to capture highly elliptical sources, arising from their implicit regularization that favors more spherical source morphologies. However, Delaunay still outperforms elliptical shapelets in robustness against spurious subhalos and agreement in parameters across filters.
4.3 Source Regularization
Shapelet source modeling is inherently regularized by the forms of the shapelet basis functions, which implicitly constrain the types of light distributions we can capture. For example, shapelets perform poorly in capturing highly peaked light distributions, thereby imposing a level of smoothness in the source. The spurious subhalo detection with shapelets may be a result of this inherent source regularization, as a spurious subhalo may compensate for the limited source complexity imposed by regularization.
To test whether spurious subhalo detections can result from source regularization, we repeat our substructure search with regularized Delaunay source modeling, which we show in Figures 17 and 18. To enforce some level of regularization in the source, we place linear priors of on the regularization strength parameter. When searching for a subhalo with Narrow Priors on the position, we find that all substructure fits are disfavored by the BIC. Interestingly, however, a wide substructure search reveals a subhalo with in the F200W filter. We conclude that gradient regularization can indeed lead to spurious subhalo detections. This finding is consistent with Nightingale et al. (2018) and Galan et al. (2024), who demonstrate that constant source regularization, where the regularization strength is controlled by a single hyperparameter, can lead to poor reconstructions and biased parameters. They also show that such problems can be resolved with an adaptive regularization scheme (Nightingale et al. 2018).
4.4 Comparison to Global Model
In this Section, we compare the CAB parameters obtained from our local analysis to the latest global cluster lens model of Abell S1063 by Beauchesne et al. (2024). We show the best-fit global convergence and shear maps published by the authors in Figure 19, with the positions of the system 4 images indicated with white squares. The cluster lens model is made up of 250 components, including two main dark matter halos near the brightest cluster galaxy (DM-BCG) and a high concentration of galaxies in the North-East (DM-NE), and three main gas components, all modeled as dual Pseudo-Isothermal Elliptical (dPIE) profiles (Limousin et al. 2005).
We present a comparison of our local CAB parameters with the CAB parameters predicted by the global model in Figure 20. Due to a degeneracy to a mass sheet transform, only relative stretch parameters can be constrained by a local lensing analysis with CAB. Hence, we derive the stretch parameters from the global model relative to the first radial stretch, , as constrained by our local analysis. The green posteriors show the CAB stretches derived from our analysis of F115W data with Delaunay source modeling. The lilac contours show the and confidence intervals of the CAB parameters derived from the global model. The errors in the 28 sampled global model parameters of the five most massive components were propagated using Monte Carlo simulations of 1,000 random convergence and shear maps. For each map in the sample, we derived the CAB stretches by calculating the average eigenvalues of the lensing Jacobian in cutouts centered at the system 4 images. We note that this procedure underestimates the errors in the CAB parameters derived from the global model, as it does not account for uncertainties in the velocity dispersion measurements of the individual cluster member galaxies.
We find good agreement between the global and local models in all but one of the CAB stretch parameters. Tension between the global and local models is expected, as the former does not incorporate pixel-level information and may be inaccurate at predicting local lensing distortions. At the same time, CAB parameters from local analysis are generally better constrained, even when underestimating the errors in CAB parameters from the global model. These results show that incorporating constraints from CAB in global lens modeling can improve the accuracy of state-of-the-art global cluster lens models. Recent work has made significant progress in incorporating pixel-level information in global cluster models, either by directly modeling bright lensed arcs (Acebron et al. 2024, Broadhurst et al. 2025) or by calculating higher-order moments of multiple images from pixel values without performing additional lens modeling (Pignataro et al. 2021, Bergamini et al. 2021, Diego et al. 2022, Sharon et al. 2022). Summary statistics from CAB analysis can complement these efforts, providing valuable information for global cluster models while maintaining the required computational efficiency to make the problem tractable.
5 Discussion & Conclusions
The CAB formalism is a promising method for probing dark matter substructure in strong cluster lenses. This work presents the first multi-filter study with CAB on real data to empirically test the method’s robustness. We find that our results are highly dependent on the source modeling method. Shapelet source modeling leads to significant systematics that can bias the CAB parameters. In contrast, pixel-based source reconstruction with Delaunay triangulation is a more robust method that suffers from significantly less systematics, as seen in the better agreement of the CAB parameters across filters.
At the same time, CAB is a local formalism that should be used with caution. Overestimating the region of validity can lead to the failure of the CAB formalism, leading to lens modeling systematics that similarly bias the inferred parameters. A multi-band analysis in different mask sizes allows us to detect the failure of the local lensing formalism as a tension between lens parameters across filters. We demonstrate that CAB works better in smaller masks for all source reconstruction methods, as seen in reduced tension between parameters from different filters. Hence, we show the power of multi-band analysis as an empirical approach to determine the region of validity of this local formalism.
Source and lens modeling systematics have profound implications for substructure search. Our statistically significant detection of a subhalo with Cartesian shapelets tells a cautionary tale that source modeling systematics can lead to convincing substructure detections. Switching to elliptical shapelets weakens the statistical preference for the spurious subhalo. This suggests that insufficient complexity of the source model can lead to a preference for a subhalo that accounts for some of the missing complexity in the source-light distribution. At the same time, the higher level of lens modeling systematics seen in Large Masks exacerbates spurious detections. As with the source model, a subhalo can likewise account for some of the missing complexity in the lens model. Multi-band analysis is key to disentangling source and lens modeling systematics from substructure detection. While seemingly convincing, several qualities give away hints that the perturber detected with shapelets may be spurious.
The smooth CAB parameters obtained with shapelet source modeling across different filters are in tension, putting into question the validity of the source and the lens models.
Some subhalo parameters are in tension across filters when using a shapelet source model. In particular, we find an anomalously large concentration in both mask sizes in the F150W filter.
The statistical significance of the subhalo detection diminishes as we complicate the source model by increasing the shapelet order parameter or allowing for ellipticity of the shapelet basis.
By introducing pixel-based source modeling with Delaunay triangulation, we allow for a highly flexible source model without any explicit regularization. With this higher source complexity, the preference of our model for substructure disappears. Interestingly, when imposing gradient regularization on Delaunay source modeling, spurious substructure appears again, although at different location, in two filters. This finding demonstrates that strong, non-adaptive regularization conditions can lead to spurious substructure detections, as additional complexity in the lens model accounts for the limited complexity in source modeling. This result is particularly relevant for JWST cluster lenses, where multiply-lensed, magnified images offer remarkably high resolutions, and a highly detailed source model becomes essential. However, we expect the result to translate to galaxy-galaxy strong lenses, where strong regularization conditions that limit the source model complexity could lead to a spurious preference for substructure that accounts for some of this missing complexity. The availability of multi-filter data with JWST gives us a higher chance of decisively ruling out spurious detections.
Finally, our local CAB model is in good agreement with the latest global model of Abell S1063. While this speaks for the success of novel global lens modeling techniques, we still place more trust in the ability of CAB to predict local deflections, as global models typically neglect pixel-level information due to computational challenges. To this end, summary statistics extracted from a CAB analysis can be incorporated to enhance the accuracy of global cluster models while keeping the problem computationally tractable.
Acknowledgements
We thank Arthur Tsang for helpful discussions. CD is partially supported by the Department of Energy (DOE) Grant No. DE-SC0020223. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program .
Data Availability
The data is publicly available on the Mikulski Archive for Space Telescopes (MAST).
Appendix A Posteriors from Large Masks
We show the full posterior distribution of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling methods, using Large Masks, in Figures 21 and 22.
Appendix B Posteriors from Small Masks
We show the full posterior distribution of the CAB parameters obtained with Cartesian shapelet and Delaunay source modeling methods, using Small Masks, in Figures 23 and 24. We compare the means and errors from the two source modeling methods in Figure 25.
Appendix C Spurious subhalo statistics
We show the BIC optimization curves of the smooth CAB and substructure models in Figures 26, 27, and 28, as well as the posterior distributions of the subhalo parameters at each shapelet order parameter in Figures 29, 30, and 31. Notably, the subhalo parameters are consistent across a broad range of shapelet order parameters, which shows that even more complex shapelet source models can suffer from spurious perturber detections.
Appendix D Elliptical Shapelets
We repeat the smooth CAB and substructure analyses using an elliptical instead of Cartesian shapelet basis. We show the resulting posterior distributions from CAB analyses in Large and Small Masks in Figures 32 and 33, respectively. We show the BIC statistic of smooth and substructure CAB models for a range of shapelet-order parameters in Figures 34, 35, and 36. We compare the statistical significance of subhalo detection between analyses relying on Cartesian and elliptical shapelet bases for source reconstruction in Figures 37 and 38.
Appendix E Mock Subhalo Tests
To compare the performance of shapelet and Delaunay source modeling methods in detecting substructures in real data, we test subhalo recovery in realistic mocks of JWST quality. We lens a real high-resolution source NGC 1300, a barred spiral galaxy, with smooth CAB deflections and a single subhalo in the line of sight of one of the lensed images. We choose a subhalo mass of , concentration of , position of (, ) = (", ") relative to the brightest lensed image pixel, at the same subhalo redshift of as the Abell S1063 cluster. The brightness, sizes, noise levels, and Point Spread Functions of the mocks are identical to our JWST observations of system 4 in filter F115W.
We fit the mock images with Cartesian shapelet and Delaunay source modeling methods, using the same masks. In shapelet source modeling, we minimized the BIC of the residuals using the true CAB and NFW parameters to find the optimal shapelet order . In Delaunay source modeling, we pick half of the unmasked pixels in the first mock image as candidate tesselation points, as described in Section 3.2.2. If any of the chosen candidate points end up being two neighboring pixels, one of them is removed.
We show our mock images, model reconstructions, and residuals from Delaunay and Cartesian shapelet source modeling in Figure 39, as well as the posterior distribution of the CAB and NFW parameters in Figure 40. We find that Delaunay is consistently less biased than Cartesian shapelets in measuring the smooth CAB parameters, which contributes to a more accurate measurement of the subhalo mass and concentration. We conclude that Delaunay is a reliable source modeling method for a robust substructure search, especially in lensing systems involving source galaxies with complex morphologies.